OpenStudy (hba):

if f(x)=root(36-x^2) then the range of values of f is given by ?

OpenStudy (hba):

@waterineyes Will i have to put root(36-x^2)not equal to zero ?

OpenStudy (anonymous):

This time you have to find range.. For that put f(x) = y Now : \[y = \sqrt{36 - x^2}\]

OpenStudy (hba):

i got this how to solve it ?

OpenStudy (anonymous):

Now convert this in terms of y : \[y^2 = 36 - x^2 \implies x^2 = 36 - y^2 \implies x = \sqrt{36-y^2}\]

OpenStudy (hba):

I got this too,Next ?

OpenStudy (anonymous):

Now put : \(36-y^2 \ne 0\) And now find y..

OpenStudy (hba):

y is not equal to plus minus 6

OpenStudy (hba):

{y:0<=y<=36} < Can this be the answer ?

OpenStudy (hba):

The other options are, {y:-6<=y<=6} {y:0<=y<=6} {y:-6<=y<=0}

OpenStudy (anonymous):

36 or 6??

OpenStudy (hba):

@waterineyes Yes ?

OpenStudy (hba):

The first one

OpenStudy (hba):

{y:0<=y<=36}

OpenStudy (hba):

Should be the answer ^

OpenStudy (hba):

@waterineyes Please help :'(

OpenStudy (hba):

@waterineyes What ?

OpenStudy (anonymous):

See from here : \[y = \sqrt{36-x^2}\] y cannot take negative values, so y cannot be : -6, -5, -4 , -3, -2, -1...

OpenStudy (anonymous):

So I think \(0 \le y \le 6\)..

OpenStudy (hba):

But y is not equal to +-6

OpenStudy (hba):

@waterineyes ?

OpenStudy (anonymous):

Sorry, I may be wrong.. I don't know..

OpenStudy (hba):

@campbell_st Please help ?

OpenStudy (hba):

@waterineyes Thanks Alot for your help.

OpenStudy (anonymous):

Gotta go.. Good Night hba..

OpenStudy (hba):

@waterineyes Good night :)

OpenStudy (campbell_st):

ok.... you can't take the square root of a negative number...does that make sense...?

OpenStudy (hba):

No.

OpenStudy (anonymous):

\[36 -x^2 \ge 0\]

OpenStudy (hba):

I am done with this :(

OpenStudy (hba):

Because no one is helping me out :(

OpenStudy (campbell_st):

start with the basics.... can you take the square root of a negative number...?

OpenStudy (hba):

Yes.

OpenStudy (hba):

It is just gonna turn complex

OpenStudy (campbell_st):

ok.... what in \[\sqrt{-4} = ?\]

OpenStudy (hba):

so 2i

OpenStudy (hba):

I mean +- 2i

OpenStudy (campbell_st):

ok... so from what you are saying you can input (the domain) any value of x....is that correct...

OpenStudy (campbell_st):

then you need to think about what output(the range) you will get...

OpenStudy (hba):

Yes.

OpenStudy (hba):

Ok so I have worked it out and the range y should not be equal to +-6

OpenStudy (campbell_st):

but the way I read this question, you are asked for the range, based on the set of real numbers...

OpenStudy (campbell_st):

no... what you have worked out with + - 6 is the input..... from the set of real numbers.... now you have to find the output... start with the largest input value 6 what is the output..?

OpenStudy (campbell_st):

positives and negatives squared give the same answers

OpenStudy (hba):

0

OpenStudy (hba):

For 6

OpenStudy (campbell_st):

ok... so input the smallest value 0... whats the output...

OpenStudy (hba):

36 ?

OpenStudy (hba):

@campbell_st ?

OpenStudy (campbell_st):

umm nope isn't it \[\sqrt{36 - 0} = 6\]

OpenStudy (campbell_st):

so the output ranges from 0 to 6

OpenStudy (hba):

Oh Yeah sorry its 6

OpenStudy (hba):

So C is the answer ?

OpenStudy (campbell_st):

you didn't post a list of your answer choices.... but basically the range(output) is determined by the domain(input)

OpenStudy (hba):

{y:0<=y<=6}

OpenStudy (hba):

@campbell_st Is this correct ^^ ?

OpenStudy (campbell_st):

that looks like it

OpenStudy (hba):

thanks :D